We select at a subject who had mammogram.find the probability that she is divorced or separated?


1
Expert's answer
2021-06-21T10:46:20-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
 Marital & Hispanic & Non- & Total \\ 
Status & & Hispanic & \\
 \hline
Currently & 319 & 738 & 1057\\
Married & & & \\
\hdashline
Divorced\ or & 130 & 329 & 459\\
Separated & & & \\
\hdashline
Widowed & 88 & 402 & 490\\
 & & & \\
\hdashline
Never\ Married & & & \\
or\ Living\ As & 41 & 95 & 136 \\
an\ Unmarried\ Couple & & & \\
\hdashline
 Total & 578 & 1564 & 2142\\
\hline
\end{array}"



"P(Divorced\ or\ Separated)=\dfrac{459}{2142}"

"=\dfrac{3}{14}\approx0.2143"